OpenStudy (anonymous):

What is 8.3 repeating as a fraction ?

OpenStudy (aum):

Let x = 8.33333.... 10x = 83.33333... Subtract first from the second: 10x - x = 83.33333... - 8.33333.... 9x = 75 x = 75/9 = 25/3